Acme Tobacco is currently selling 5,000 pounds of pipe tobacco per year. Due to competitive pressures, the average price of a pipe declines from $15 to $12. As a result , the demand for Acme pipe tobacco increases to 6,000 pounds per year.


a) what is the cross elasticity of demand for pipes and pipe tobacco?


b) Assuming that the cross elasticity does not change, at what price of pipes would the demand for pipe tobacco be 3,000 pounds per year? Use $15 as the initial price of a pipe.



1
Expert's answer
2022-03-18T12:00:17-0400

a) Cross Elasticity= "\frac{\Delta Q_A}{\Delta P_p}\times \frac{P_p}{Q_A}"

"\frac{6000-5000}{12-15}\times \frac{15}{5000}"

"\frac{1000}{-3}\times \frac{15}{5000}= -1"


b) "\frac{3000-5000}{\Delta P}\times \frac{15}{5000}"

"=\frac{-2000}{\Delta P }\times \frac{15}{5000}"

"\Delta P= \frac{-2000 \times 15}{5000}= -30"

Price = 30+15= 45